OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a 12 -inch diameter pizza and an 18 -inch diameter pizza. He has set the price of a 12 -inch cheese pizza at $8.00 , based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an 18-inch cheese pizza? Explain .

OpenStudy (cwtan):

Using ratio pi 144:pi 324=8:x 144/324=8/x x=8(324/144) =18 I am not sure about this

OpenStudy (cwtan):

oops i misused the diameter as radius

OpenStudy (anonymous):

so how can i fix it ?

OpenStudy (cwtan):

using ratio Area of 12-inch pizza: Area of 18-inch pizza=Price of 12-inch pizza:Price of 18-inch pizza pi 36:pi 81=8:x I removed pi for easier calculation 36/81=8/x x=8(81/36)=18 lol same answer

OpenStudy (anonymous):

woww , thank you (: